OpenStudy (jjuden):

If f(x) varies directly with x and f(x) = 50 when x = –10, then what is f(x) when x = 15?

OpenStudy (jjuden):

–150 –75 –30 –5

OpenStudy (mathmale):

Mind sticking to helping jjuden with her valid question? Please?

OpenStudy (mathmale):

@Thatguy12310 : You've been warned: there will be negative consequences for your abuse of this learning site.

OpenStudy (blues):

Thatguy, please don't spam other peoples' questions. :)

OpenStudy (mathmale):

If f(x) varies directly with x and f(x) = 50 when x = –10, then what is f(x) when x = 15? I'd suggest you use this model for direct variation f(x) = kx Now substitute the givens: f(x) = 50 and x = -10. Calculate the constant of proportion, k.

OpenStudy (jjuden):

i dont follow @mathmale

OpenStudy (anonymous):

If f(x) varies directly with x and f(x) = 50 when x = –10, then what is f(x) when x = 15? not sure but maybe this c x =f(x) so c * -10 =50 solve for c c=50/-10 = -5 so -5*x=f(x) now f(15)=-5*15 = -75 se how i did it.

OpenStudy (jjuden):

yea @timo86m

OpenStudy (mathmale):

jjuden: What do you know about "direct proportation"? This problem does not mention "direct proportion," but that is the underlying concept here. You need to find the constant of proportionality, k, which Tim calls "c" (that's perfectly OK). Once you have this, then you have a relationship between f(x) and x. Using that relationship, you calculate f(x) for x =15.

OpenStudy (jjuden):

i got -5 is that right i did the math twice